# Cost-Optimal Grouped-Query Attention for Long-Context Modeling

### v2

@@ Abstract
Grouped-Query Attention (GQA) is a widely adopted strategy for reducing the computational cost of attention layers in large language models (LLMs). However, current GQA configurations are often suboptimal because they overlook how context length influences inference cost. Since inference cost grows with context length, the most cost-efficient GQA configuration should also vary accordingly. In this work, we analyze the relationship among context length, model size, GQA configuration, and model loss, and introduce two innovations: (1) we decouple the total head size from the hidden size, enabling more flexible control over attention FLOPs; and (2) we jointly optimize the model size and the GQA configuration to arrive at a better allocation of inference resources between attention layers and other components. Our analysis reveals that commonly used GQA configurations are highly suboptimal for long-context scenarios. More importantly, we propose a recipe for deriving cost-optimal GQA configurations. Our results show that for long-context scenarios, one should use fewer attention heads while scaling up model size. Configurations selected by our recipe can reduce both memory usage and FLOPs by more than 50% compared to Llama-3’s GQA, with no degradation in model capabilities. Our findings offer valuable insights for designing efficient long-context LLMs. [^1] The code is available at: https://www.github.com/THUNLP/cost-optimal-gqa
@@
Yingfa Chen1*, Yutong Wu2*, Chenyang Song1, Zhen Leng Thai1, Xingyu Shen1, Xu Han1, Zhiyuan Liu1, Maosong Sun1 1NLP Group, DCST, IAI, BNRIST, Tsinghua University, Beijing, China 2SIST, University of Science and Technology Beijing, Beijing, China chenyingfa1999@gmail.com, wuyutong_yuna@163.com
footnotetext: Equal contributions
@@ 1 Introduction
One of the most widely used techniques for reducing memory costs is Grouped-Query Attention (GQA) (Ainslie et al., 2023), in which attention heads are split into groups and the heads in each group share the same KV vectors. Current implementations of GQA have two critical limitations: (1) Most existing models unnecessarily restrict the total number of head dimensions to be equal to the model hidden size, resulting in redundant FLOPs (floating-point operations). (2) When deciding on the number of attention heads and groups, current models do not take into account the influence of context length on the computational and memory costs, resulting in suboptimal configurations for long contexts.
@@ 2 Related Work
$N$ Model size None
@@ 4 Method
Our objective is to find the GQA configuration that minimizes inference costs while attaining a given loss. We approach this by framing the problem as balancing the time-variant and time-invariant costs.
In order to unlock the ability to flexibly allocate different amounts of compute and memory to the time-variant and time-invariant components, we make two changes to existing GQA design procedures: (1) We decouple the number of attention heads from the model hidden dimension, and (2) we jointly optimize the model size and the GQA configuration. Figure 1 (left) shows the effect of these two changes, and Table 1 shows the adjustability of different hyperparameters in this work compared to vanilla GQA.
@@ 4.1 Cost-Optimal GQA Search
$H_{\text{cand}}$ =$=\{n_{h}=1,2,4,\dots,\max(d)/d_{h}\}$ (6)
×$\,\,\,\,\times\{n_{kv}=1,2,4,\dots,\max(d)/d_{h}\}$
[…]
@@ 5.1 Experimental Settings
We adopt the widely used Llama-3 (Grattafiori et al., 2024) architecture. For each GQA configuration, we train models from 3M to 1.2B in size. We keep the model configurations as close to Biderman et al. (2023) as possible. We have $\max(d)/d_{h}=32$, this results in 21 candidate configurations (i.e., $|H_{\text{cand}}|=21$). For more details, see Appendix D.
@@
We use SlimPajama (Soboleva et al., 2023) in our experiments. It is a deduplicated version of the RedPajama (Weber et al., 2024) corpus with 627B tokens. In most of our experiments, we use a 20:1 ratio between training data and model parameters, as suggested by Hoffmann et al. (2022). Additionally, we always ensure that each batch has 512K tokens. For more details, see Appendix E.
@@
We try to follow common practices in most of our experiments. We use AdamW optimizer with the WSD learning rate scheduler (Hu et al., 2024). We choose the maximum learning rate by sweeping different values with the MHA model for each model size. For more details, see Appendix F.
@@
We discover that loss does not have a simple relationship (e.g., power-plus-constant function) with either memory or computational costs. However, it is still possible to predict the loss by fitting loss as a function of $N$, then transforming the fitted curves along the x-axis to account for the time-variant costs. Fitting loss as a power-plus-constant function of $N$ is highly accurate, with $R^{2}$ values over 0.999.
@@ 2.35 $32,32$ $32,32$ $32,32$ $32,16$ $32,8$
Table 4 reports the cost-optimal GQA for different expected inference context lengths $T$ and target losses $\mathcal{L}^{*}$. When the target loss is high, the model is small, making the time-invariant costs low. Thus, the optimal configuration allocates more resources to the time-invariant part by increasing $N$ and reducing $n_{h}$ and $n_{kv}$. Similarly, when $T$ is great, the time-variant costs are high, making it more attractive to reduce $n_{h}$ and $n_{kv}$ more aggressively.
In addition, the results indicate that there is nothing especially attractive about the commonly used Llama-3 GQA configuration ($d/d_{h},8$). For certain combinations of $\mathcal{L}^{*}$ and $T$, the GQA configuration is cost-optimal. However, for a greater number of combinations, it is sub-optimal in costs. The result implies that configuration of GQA Transformers should take into account the expected inference context length, and directly applying the popular GQA configuration results in severe waste of hardware resources.
@@ 5.6 Aligning Training Costs
Figure 4 reports the result when we always train with $T=128$K [^6] LMs are usually trained with short contexts most of the time, so this result may not apply.. We find that using fewer heads is even more advantageous because of the extra training data, producing a model with the same loss but with 88% and 83% lower memory and FLOPs usage.
@@ 5.7 Influence of Context Length
Figure 5 shows the relative loss difference between various GQA configurations with $H=1,1$ as the baseline. Figure 6 shows this relationship when varying $N$, with $N$=150M as the baseline. The results show that the relative loss difference is relatively flat when $T>$ 8K (all fluctuations are less than 1%). The main takeaway is that when applying our cost optimization procedure to longer contexts, we do not have to repeat step 2 (an expensive process) with longer contexts since the loss change of each model will remain roughly the same.
@@ 6 Conclusion
To optimize the allocation of FLOPs and memory between time-invariant and time-variant components of GQA Transformers, we first decouple the number of attention heads from the model hidden dimensions, enabling a more flexible distribution of FLOPs and memory. Next, we refine the estimation of computational and memory costs in existing approaches by incorporating context length. Our findings reveal that typical configurations of GQA are significantly suboptimal for certain context lengths. Through detailed analysis, we offer valuable insights for improving the allocation of resources by jointly adjusting the model size and the number of query and KV heads. As the demand for greater inference context lengths continues to grow, our work marks a critical advancement toward efficient long-context LLMs.
@@
Like most phenomena in neural language models, we cannot be sure that the conclusions will hold when further scaling up the models. The power-plus-constant scaling law is also not guaranteed, although it has been empirically validated up to hundreds of billions of parameters. Similarly, there is no guarantee that these laws and our conclusions will hold for an arbitrarily large amount of training data. In general, we have kept our experiments as close to research conventions as possible, and the scale of the largest models in our experiments (i.e., 1.2B for Llama-3 GQA and 1.8B for our cost-optimal GQA) is comparable to some real-world LLMs.
@@ Appendix A Notations
$N$ Number of model parameters.
@@ Appendix C Training Costs of GQA Transformers
Infer. Mem. $(M_{\text{infer}})$ $N$ $2TLd_{h}n_{kv}$
@@
Train. Mem. $(M_{\text{train}})$ $4N$ $TdL$
@@ I.1 Constant Number of KV Heads
This section provides additional details for Section 5.5.

### v3

@@ Abstract
Grouped-Query Attention (GQA) is a widely adopted strategy for reducing the computational cost of attention layers in large language models (LLMs). However, current GQA configurations are often suboptimal because they overlook how context length influences inference cost. Since inference cost grows with context length, the most cost-efficient GQA configuration should vary accordingly. In this work, we analyze the relationship among context length, model size, GQA configuration, and model loss, and introduce two innovations: (1) we decouple the total head size from the hidden size, enabling more flexible control over attention FLOPs; and (2) we jointly optimize the model size and the GQA configuration to arrive at a better allocation of inference resources between attention layers and other components. Our analysis reveals that commonly used GQA configurations are highly suboptimal for long-context scenarios. Moreover, we propose a recipe for deriving cost-optimal GQA configurations. Our results show that for long-context scenarios, one should use fewer attention heads while scaling up the model size. Configurations selected by our recipe can reduce both memory usage and FLOPs by more than 50% compared to Llama-3’s GQA, with no degradation in model capabilities. Our findings offer valuable insights for designing efficient long-context LLMs. [^1] The code and models are available at https://www.github.com/THUNLP/cost-optimal-gqa.
@@
Yingfa Chen1∗, Yutong Wu2∗, Chenyang Song1, Zhenleng Thai1, Xingyu Shen1, Xu Han1†, Zhiyuan Liu1†,  and Maosong Sun1 1NLP Group, DCST, IAI, BNRIST, Tsinghua University, Beijing, China 2SIST, University of Science and Technology Beijing, Beijing, China chenyingfa1999@gmail.com, wuyutong_yuna@163.com {han-xu,liuzy}@tsinghua.edu.cn
11footnotetext: Equal contributions.22footnotetext: Corresponding authors.
@@ 1 Introduction
One of the most widely used techniques for reducing memory costs is Grouped-Query Attention (GQA) (Ainslie et al., 2023), in which attention heads are split into groups and the heads in each group share the same KV vectors. Current implementations of GQA have two critical limitations: (1) Most existing models unnecessarily restrict the total number of head dimensions to be equal to the hidden size, resulting in redundant FLOPs (floating-point operations). (2) When deciding on the number of attention heads and groups, current models do not take into account the influence of context length on the computational and memory costs, resulting in suboptimal long-context configurations.
@@ 2 Related Work
NN Model size None
@@ 4 Method
Our objective is to find the GQA configuration that minimizes inference costs while attaining a given loss. We approach this by framing the problem as balancing the time-variant and time-invariant costs. In order to unlock the ability to flexibly allocate different amounts of compute and memory to the time-variant and time-invariant components, we make two changes to existing GQA design procedures: (1) We decouple the number of attention heads from the model hidden dimension, and (2) we jointly optimize the model size and the GQA configuration. Figure 1 (left) shows the effect of these two changes, and Table 1 shows the adjustability of different hyperparameters in this work compared to vanilla GQA.
@@ 4.1 Cost-Optimal GQA Search
$H_{\text{cand}}$ $=\{n_{h}=1,2,4,\dots,\max(d)/d_{h}\}$ (6)
$\,\,\,\,\times\{n_{kv}=1,2,4,\dots,\max(d)/d_{h}\}$
$\text{s.t. }n_{kv}\leq n_{h},$
@@ 5.1 Experimental Settings
More details of the experimental settings are in Appendix D, Appendix E, and Appendix F.
@@
We adopt the popular Llama-3 (Grattafiori et al., 2024) architecture. For each GQA configuration, we train models from 3M to 1.2B in size. We keep the model configurations as close as possible to Biderman et al. (2023). We have $\max(d)/d_{h}=32$, this results in 21 candidate configurations (i.e., $|H_{\text{cand}}|=21$).
@@
We use SlimPajama (Soboleva et al., 2023) in our experiments. It is a deduplicated version of the RedPajama corpus (Weber et al., 2024) with 627B tokens. In most of our experiments, we use a 20:1 ratio between training data and model parameters, as suggested by Hoffmann et al. (2022). Additionally, we always ensure that each batch has 512K tokens.
@@
We follow common practices in most of our experiments. We use the AdamW optimizer with the WSD learning rate scheduler (Hu et al., 2024). We choose the maximum learning rate by sweeping different values with the MHA model for each model size.
@@
We find that loss does not have a simple relationship (e.g., power-plus-constant function) with either memory or computational costs. However, it is possible to predict the loss by fitting the loss as a function of $N$, then transforming the fitted curves along the x-axis to account for the time-variant costs. Fitting loss as a power-plus-constant function of $N$ is highly accurate, with $R^{2}$ values over 0.999.
@@ 2.35 $32,32$ $32,32$ $32,32$ $32,16$ $32,8$
Table 4 reports the cost-optimal GQA for different expected inference context lengths $T$ and target losses $\mathcal{L}^{*}$. When the target loss is high, the model is small, making the time-invariant costs low. Thus, the optimal configuration allocates more resources to the time-invariant part by increasing $N$ and reducing $n_{h}$ and $n_{kv}$. Similarly, when $T$ is great, the time-variant costs are high, making it more attractive to reduce $n_{h}$ and $n_{kv}$ more aggressively. The results also indicate that there is nothing especially attractive about the commonly used Llama-3 GQA configuration ($d/d_{h},8$). For certain combinations of $\mathcal{L}^{*}$ and $T$, the GQA configuration is cost-optimal. However, for a greater number of combinations, it is sub-optimal. The result implies that the configuration of GQA Transformers should consider the expected inference context length. Directly applying the popular GQA configuration results in wasting hardware resources.
@@ 5.6 Aligning Training Costs
Figure 4 reports the result when we always train with T=128T=128K [^6] LMs are usually trained with short contexts most of the time, so this result may not apply.. We find that using fewer heads is even more advantageous due to the additional training data, resulting in a model with the same loss but with 88% and 83% lower memory and FLOPs usage.
@@ 5.7 Influence of Context Length
Figure 5 shows the relative loss difference between various GQA configurations with $H=1,1$ as the baseline. Figure 6 shows this relationship when varying $N$, with NN=150M as the baseline. The results show that the relative loss difference is relatively flat when $T>$ 8K (all fluctuations are less than 1%). The main takeaway is that when applying our cost optimization procedure to longer contexts, we do not have to repeat step 2 (an expensive process) with longer contexts since the loss change of each model will remain roughly the same.
@@ 6 Conclusion
To optimize the allocation of FLOPs and memory between time-invariant and time-variant components of GQA Transformers, we first decouple the number of attention heads from the model hidden dimensions, enabling a more flexible distribution of FLOPs and memory. Next, we refine the estimation of computational and memory costs in existing approaches by incorporating context length. Our findings reveal that typical configurations of GQA are significantly suboptimal for specific context lengths. Through detailed analysis, we provide valuable insights for improving resource allocation by jointly adjusting the model size and the number of attention heads. As the demand for greater inference context lengths continues to grow, our work marks a critical advancement toward efficient long-context LLMs.
@@
Like most phenomena in neural language models, we cannot be certain that the conclusions will hold when the models are further scaled up. The power-plus-constant scaling law is also not guaranteed, although it has been empirically validated up to hundreds of billions of parameters. Similarly, there is no guarantee that these laws and our conclusions will hold for an arbitrarily large amount of training data. In general, we have kept our experiments close to research conventions, and the scale of the largest models in our experiments (i.e., 1.2B for Llama-3 GQA and 1.8B for our cost-optimal GQA) is comparable to some real-world LLMs.
We have not thoroughly ablated the influence of all possible hyperparameters due to limited resources. Some important hyperparameters that may affect our conclusions include the head dimension, vocabulary size, and model aspect ratio. More extensive ablation studies are a promising future research direction.
More recent LLMs have adopted even more advanced techniques that complicate the calculations of inference costs. Some notable techniques include speculative decoding, sparse mixture-of-experts, and hybrid recurrence-attention models. Applying our approach to such models remains a promising area for future research.
Acknowledgments
This work is supported by the National Key Research and Development Program of China (2024YFB4505603) and a grant from the Guoqiang Institute, Tsinghua University.
@@ Appendix A Notations
NN Number of model parameters.
@@ Appendix C Training Costs of GQA Transformers
Infer. Mem. $(M_{\text{infer}})$ NN $2TLd_{h}n_{kv}$
@@
Train. Mem. $(M_{\text{train}})$ $4N$ TdLTdL
@@ I.1 Constant Number of KV Heads
This section provides details for Section 5.5.